For the series below, (a) find the series radius and interval of convergence. For what values of x does the series convergeFor the series below, (a) find the series radius and interval of convergence. For what values of x does the series converge

i need show work for this problem

pls clear show work

cal 2
Answer #1

Given series 3 nok? to find isterial Applejing series ratio test of Covergence. La chimenea Conti nto aan = nu nanti = (n+8th 9t! in ntos I I (MA)ht x H nox? ntl nta nx. La din | M41) neoten (a) din flas. din gra) uta xta ta lin 1 (ot) too l 1. din l into Ial. 00 1 (10(146) 000lim (1++)?= ese. n0 490 = x(os) ( din fitoh) din (1462) 500.09 C (11. e) = a(e)s a IXI AS L= ∞(x) = 0AS So No L=0 By Ratio fest the series dwerges interval of covergens Interwal of convergence is X=0 Radies of Convergence is a(a) The convergence series radies of y lol series isterial of The Covergence is A N=0 (6) A. X=0 Cc (D. the series Conditiona